### v1

@@ 3.2.1 Rotation of lines
Figure 4: A visualization of one of the experimental results. The blue line shows the output of the QNN for the input test data (red line) after training. The purple line shows the output of the RQNN for the input test data (red line) after training.
Table 4: Average and standard deviation of the angles between the QNN output and the RQNN output over 50 trials
Subject Avg $\pm$ std
Angle between QNN output and RQNN output 102.5 $\pm$ 48.0
Angle between QNN output and input test data 94.4 $\pm$ 29.9
Angle between RQNN output and input test data 78.9 $\pm$ 39.8
Table 5: Euler angles of rotation from the input test data to the QNN and the RQNN outputs for the result shown in Fig. 4.
i (roll) j (pitch) k (yaw)
Input test data – QNN output $-76.5$ $14.1$ $-11.1$
Input test data – RQNN output $-23.9$ $-77.7$ $19.3$
3.2.2 Rotation of point cloud
Subsequently, an experiment on rotational representation was conducted using the plane point cloud data shown in the left figure in Fig. 5. These training data consisted of a rectangle composed of 5,000 points on the ij-plane, both vertically and horizontally ranging from 0 to 1, which was used as the input data. The plane rotated 90 degrees around the j-axis was used as the teacher data.
Both the QNN and the RQNN constructed a four-layer fully connected neural network with a 1-25-50-1 architecture. In this experiment as well, since the training data involved linear transformations, an identity function was used as the activation function for all neurons in each layer instead of a nonlinear function. The learning rate was set to 0.05. The Adam optimizer was used, and learning was conducted 10 times until the loss value was less than 0.01. Additionally, the weights were initialized using the Xavier uniform distribution, and the initial values for the RQNN and the QNN were unified.
Figure 5: The left figure shows the training data used in the point cloud rotation experiment. The blue points represent the input data, and the orange points represent the output data. The right figure depicts the test data consisting of 5,000 points forming a quadrilateral pyramid.
As the test data, a quadrilateral pyramid consisting of 5,000 points, shown in the right figure of Fig. 5, was used. Fig. 6 visualizes the output values of the QNN and the RQNN for the input test data in one of the ten trials conducted. It can be seen from Fig. 6 that the output of the RQNN is different from the output of the QNN for the same input test data. To quantitatively evaluate the overall motion of this quadrilateral pyramid, we chose the four representative points (red, blue, yellow, and green) characterizing the features of the quadrilateral pyramid, and calculated the Euler angles of rotation of the four vertices of the tetrahedron. Table 6 shows the Euler angles (Eqs. (23) - (25)) of the rotation of the four vertices from the input test data to the outputs of the QNN and RQNN for the result shown in Fig. 6. It is also obvious from Table 6 that the QNN and the RQNN output completely different values for the same input test data.
For reference, the results of using a nonlinear activation function for the RQNN and QNN in this experiment are shown in Fig. 7 of the appendix. The parameters and training conditions for both models were the same, with the sigmoid function used in the hidden layers and the identity function used in the output layer. As expected, these also produced completely different outputs for the same input test data.
Figure 6: Outputs of the QNN and the RQNN for the input test data shown in Fig. 5 in the experiment on the point cloud rotation
Table 6: Euler angles of rotation of the four vertices from the input test data to the QNN and the RQNN outputs for the result shown in Fig. 6.
Input test data – QNN output
i (roll) j (pitch) k (yaw)
Blue $28.4$ $3.1$ $-12.1$
Red $-5.4$ $-8.8$ $14.1$
Green $-37.6$ $13.2$ $-4.5$
Yellow $-15.5$ $-33.3$ $48.9$
Input test data – RQNN output
i (roll) j (pitch) k (yaw)
Blue $3.8$ $1.6$ $-44.9$
Red $-1.0$ $3.0$ $-2.1$
Green $-38.1$ $34.6$ $-12.3$
Yellow $-2.8$ $-18.8$ $16.5$
4 Discussion
The learning speed of the RQNN is several times faster than the Real-NN, but there was no significant difference in learning speed compared to the traditional QNN. As seen in Table 3 and Fig. 2, despite having higher initial losses than the Real-NN, both the QNN and the RQNN converged about three times faster on average. However, between the RQNN and the QNN, the RQNN completed learning in 3,306 epochs, while the QNN took about 3,111 epochs on average, showing no significant difference. This is likely because both the RQNN and the QNN are neural network architectures based on quaternions.
The rotational generalization performance of the RQNN was observed in both the rotation of lines and the rotation of objects generated from point clouds. It is clear from Fig. 4 that there are differences in the output results between the QNN and the RQNN for the input test patterns. Furthermore, as seen in Table 4, the average angle of the outputs between the QNN and the RQNN was 102 degrees, suggesting that the RQNN learned a different rotation than the QNN. The angles between the input test data and the outputs of the QNN and the RQNN are 94.4 and 78.8 degrees on average, respectively, which are similar to the value of angle (i.e., 90 degrees) given during the training. Additionally, examining the rotation angles from the input test data to the QNN and the RQNN using the Euler angles reveals that the outputs of the two models are different (Table 5). Significant differences were observed in rotations around the j-axis, i-axis, and k-axis. Looking at the experimental result on the point cloud shown in Fig. 6, it is apparent that the attitudes of the tetrahedron and the positions of the colored vertices in both the RQNN and the QNN exhibit different behaviors. Furthermore, examining the rotation angles of the colored vertices from the test data using the Euler angles also reveals that the behavior of the RQNN differs from that of the QNN (Table 6). As a result of conducting two experiments on the generalization performance of rotation, it was discovered that the RQNN acquires a different rotational representation compared to the QNN, despite using models of the same configuration, the same dataset, and the same learning conditions.
5 Conclusion
In this paper, we proposed a new multi-layer feedforward quaternion neural network model architecture, the Reverse Quaternion Neural Network, utilizing the non-commutative nature of quaternion multiplication, and investigated its learning characteristics. Comparing the results of the experiments on convergence speed and the rotational generalization performance with the traditional models, it was found that the learning speed of the Reverse Quaternion Neural Network is similar to that of the conventional quaternion neural network, but the Reverse Quaternion Neural Network has a different generalization on rotation.
As a future challenge, it is important to further investigate the learning representation characteristics of the RQNN. In this paper, relatively simple experiments were conducted, and differences in behavior were identified, but it is necessary to explore the potential of the RQNN through experiments using more complex learning datasets. Moreover, regarding the generalization performance on rotation, it is also necessary to investigate from a theoretical perspective. By mathematical analyses, it would be clarified in which areas the RQNN has advantages compared to traditional models.
6 Appendix
Figure 7: Outputs of the QNN and the RQNN using the sigmoid function in the hidden layers and the identity function in the output layer in the point cloud rotation experiment for the input test data shown in Fig. 5

### v2

@@ III-A Rotation of Lines Task
Figure 3: Visualization of four experimental trials out of fifty trials. The blue line shows the output of the QNN for the input test data (red line) after training. The purple line shows the output of the RQNN for the input test data (red line) after training.
Figure 4: Visualization of the evaluation metrics for the QNN and the RQNN outputs over 50 trials. The evaluation metrics used are as follows: (1) angle differences between tangent vectors, (2) trajectory differences (measured by Euclidean distance and curvature), (3) phase space differences in first and second derivatives, and (4) geometric consistency measured by deviation from 90 degrees orthogonality
Metric Mean Std Min Max
Angle Differences (degrees) 87.65 38.73 15.27 159.16
Trajectory Differences 0.19 0.07 0.07 0.37
Phase Space Differences 0.0036 0.0015 0.0011 0.007
Geometric Consistency (degrees) 4.56 2.89 1.070 14.93
TABLE I: Statistical analysis of four evaluation metrics computed from Eqs.(22) - (4) over 50 trials
III-B Three-dimensional Object Rotation Task
Next, we investigated the learning characteristics of the RQNN for 3D point cloud object rotation using 3D point cloud datasets. In this experiment, we utilized the ModelNet10 dataset [17], which is a subset of ModelNet40. The ModelNet10 dataset consists of ten categories: chair, table, toilet, sofa, bed, desk, dresser, nightstand, bathtub, and monitor, and has been widely adopted as a standard benchmark for 3D object recognition.
The input data underwent random rotation, and as shown in Fig. 5, these rotated samples were used as the input learning data, while the original objects served as the output target data. This configuration enabled the model to learn the transformation from arbitrary rotated orientations to the original pose. The rotation matrices were generated using Rodrigues’ rotation formula, with rotation angles constrained to a maximum of 90 degrees.
Figure 5: Visualization of 3D point cloud data: Input learning points (left) and output target points (right). The rotated object serves as the input learning data, while the original, unrotated object serves as the output target data. This approach enables the model to learn a rotation that restores the object to its original orientation.
Using this data, we developed models incorporating the RQNN and the QNN as architectural components and investigated their learning behavior by analyzing weights, activations, and gradients after training. We constructed a model architecture incorporating the QNN and the RQNN modules within the encoder structure (Fig. 6). The model consists of three main components: an encoder for feature extraction, a global feature generator, and a decoder. The input three-dimensional data is converted to pure quaternions by setting the real component to zero and assigning the data to the imaginary components, resulting in 4-dimensional quaternion features. The encoder comprises three quaternion fully-connected layers that perform quaternion multiplication operations, sequentially transforming these features through 16, 32, and 16 dimensions. Each quaternion linear layer is followed by normalization, ReLU activation, and dropout.
The global feature generator achieves higher-order representations by flattening the encoder output ($16$ dimensions $\times$ number of input points) and processing it through two fully-connected layers ($16\times 1024\rightarrow 1024\rightarrow 512$ dimensions). To enhance model expressiveness and preserve local feature information, we implemented skip connections from the encoder’s final layer ($16$ dimensions).
The decoder concatenates global features ($512$ dimensions) with skip connection features ($16$ dimensions) to generate a $528$-dimensional feature vector, which is then processed through three real-valued fully-connected layers ($528\rightarrow 256\rightarrow 128\rightarrow 3$ dimensions) to predict the final 3D coordinates.
Figure 6: Architecture of the proposed quaternion neural network model, consisting of a quaternion encoder for feature extraction, a global feature generator for higher-order representation learning, and a decoder with skip connections for 3D coordinate prediction.
The loss function was constructed as a weighted sum of MSE loss, Chamfer distance, and geometric loss, computed according to Eqs.(27) - (30). Each loss term serves a specific purpose in the training process: the MSE loss ensures point-wise accuracy, the Chamfer distance handles global point cloud alignment by measuring bidirectional point-to-point distances, and the geometric loss preserves local structural relationships by considering neighboring point configurations. This combination allows the model to maintain both global shape consistency and local geometric details during the rotation transformation.
MSE $=\frac{1}{N}\sum_{i=1}^{N}\|x_{i}-y_{i}\|^{2},$ (27)
Chamfer $=\frac{1}{B}\sum_{b=1}^{B}\left(\frac{1}{N}\sum_{i=1}^{N}\min_{j}\|x_{i}-y_{j}\|_{2}\right$.
$\qquad\left.+\frac{1}{N}\sum_{j=1}^{N}\min_{i}\|x_{i}-y_{j}\|_{2}\right),$ (28)
Structure $=\frac{1}{NK}\sum_{i=1}^{N}\sum_{k=1}^{K}\|(x_{i}-x_{i,k})-(y_{i}-y_{i,k})\|^{2},$ (29)
Loss $=\text{MSE}+\alpha\text{Chamfer}+\beta\text{Structure}.$ (30)
where $x_{i}$ and $y_{i}$ represent predicted and target point coordinates respectively, $N$ denotes the number of points, $j\in\{1,\ldots,N\}$ is the index for target points, $B$ is the batch size, $K$ represents the number of nearest neighbors, $\alpha$ and $\beta$ are the weights for Chamfer and geometric losses respectively, and $\|\cdot\|_{2}$ denotes the L2 norm (Euclidean distance). The subscripts $i,k$ in $x_{i,k}$ and $y_{i,k}$ indicate the $k$-th nearest neighbor of the $i$-th point.
The model comprises 17.5M parameters and was trained using the Adam optimizer with a batch size of 16 and a learning rate of $1\times 10^{-6}$. The loss weights $\alpha$ and $\beta$ were set to 0.1 and 0.2 respectively, with the number of nearest neighbors $K$ set to 20. Twenty percent of the training data was reserved for validation, and the model was trained for 100 epochs. The learning curves for both models are shown in Fig. 7, and the final loss values computed using Eq. (30) for the test data are presented in Table II.
Figure 7: Learning curves for models incorporating QNN and RQNN modules. Both architectures demonstrate similar convergence characteristics, exhibiting comparable behavior throughout the training process.
Model Params Loss
QNN-modules 17.5M 0.1499
RQNN-modules 17.5M 0.1501
TABLE II: Comparison of loss values between QNN and RQNN-integrated models evaluated on test data
Using the trained models, we extracted the encoder sections composed of the QNN and the RQNN architectures and analyzed the learning behavior of the three quaternion fully-connected layers for both models on the test data. The distributional characteristics of weights, activations, and gradients were visualized through Kernel Density Estimation (KDE) plots in Fig. 8, and their statistical measures are summarized in Table III.
Figure 8: Comparative analysis of models with QNN and RQNN modules. Distributions of weights (top), activations (middle), and gradients (bottom) are visualized across three encoder layers, demonstrating distinctive patterns in parameter space exploration and learning dynamics between the two architectural approaches.
Layer Model Weight_Avg Weight_Std Activation_Avg Activation_Std Gradient_Avg Gradient_Std
1 QNN 0.094638 0.415134 0.411645 0.578605 -0.006139 0.006907
1 RQNN -0.137196 0.443018 0.398682 0.565451 -0.005593 0.013652
2 QNN 0.056367 0.270802 0.394833 0.625221 -0.000385 0.006376
2 RQNN -0.018313 0.314268 0.401765 0.599941 0.000887 0.010724
3 QNN -0.017184 0.196021 0.415738 0.568344 0.001543 0.022889
3 RQNN -0.030565 0.181208 0.395756 0.600589 0.000568 0.015029
TABLE III: Statistical distributions of weights, activations, and gradients across layers in the QNN and the RQNN
As illustrated in Fig. 7, models incorporating the QNN and the RQNN modules achieved convergence with nearly identical learning trajectories. While the test data results in Table II indicate marginally better performance for the QNN model, the difference in performance metrics between the two approaches remains negligible. In contrast, analysis of Fig. 8 and Table III reveals that while both models achieved similar performance metrics, they exhibited distinct learning characteristics in feature acquisition during training, as evidenced by their different gradient patterns and weight distributions.
IV Discussion
In the rotation of lines task, the learning behavior of the RQNN exhibited distinct differences compared to conventional models. As shown in Fig. 3, the outputs for identical test data demonstrated significant variations. Analysis of Angle Differences from Table I and Fig. 4 revealed that despite some variability, the average across 50 trials was 87.65 degrees, indicating both models successfully learned 90-degree rotations. The Geometric Consistency averaging 4.56 further corroborates the high-precision achievement of 90-degree rotation learning. The Trajectory Differences averaged 0.19, indicating a consistent disparity between the two outputs. While the Geometric Consistency showed a spike around the 48th trial, it maintained stability with an average of 4.56 degrees and a standard deviation of 2.89, demonstrating the high-precision achievement of 90-degree rotation learning. The Phase Space Differences exhibited remarkably small values, with an average of 0.0036 and a standard deviation of 0.0015, indicating that both models share similar operational characteristics in their dynamic behavior. These findings indicate that while the RQNN shows differences from the QNN in terms of output orientation and trajectory, it maintains the fundamental operational characteristics.
In the 3D object rotation task using ModelNet10, both models incorporating the QNN and the RQNN modules successfully achieved learning objectives, exhibiting remarkably similar learning curves. As shown in Table II, the loss values on test data were nearly identical, indicating comparable performance between the two models. However, when comparing the quaternion layers extracted from the trained models, significant differences in weight distributions and gradient behavior were observed between the two architectures, as illustrated in Fig. 8 and Table III. Regarding weight distributions, the RQNN weights in Layer 1 and Layer 2 showed mean values of 0.137-0.137 and 0.018-0.018, with standard deviations of $0.44$ and $0.31$ respectively, demonstrating a broader distribution compared to the QNN weights. This spread suggests that the RQNN potentially explores a wider range of features during the early stages of learning. In Layer 3, both models’ weight distributions converged to similar shapes as training progressed. The observed bias towards negative mean values in the weights can be attributed to the differences in quaternion multiplication order. The activation distributions exhibited highly similar patterns across all layers for both the QNN and the RQNN, suggesting that the order of quaternion multiplication has minimal impact on activation values. However, significant differences were observed in gradient distributions between the two models. the QNN demonstrated a tendency to efficiently direct learning towards specific directions from the early stages. In contrast, the RQNN’s gradient distributions initially showed a wider range before concentrating as training progressed. This wider distribution characteristic of the RQNN potentially offers advantages in terms of learning stability and gradient vanishing prevention, while gradually refining its feature representation in deeper layers.
V Conclusion
In this paper, we proposed a new neuron structure for multi-layer feedforward quaternion neural networks called the Reverse Quaternion Neural Network (RQNN), which leverages the non-commutative property of quaternion multiplication, and investigated its learning characteristics. Through comparative analysis with conventional models in lines rotation tasks, we found that the RQNN exhibited distinct generalization capabilities. Furthermore, our experiments with 3D object rotation tasks using ModelNet10 demonstrated that while the QNN and the RQNN architectures achieved similar performance metrics, they developed fundamentally different learning representations and strategies.
Future research directions should focus on deeper investigation of RQNN’s learning characteristics and representation properties. While our current study demonstrated unique generalization properties through fundamental experiments, comprehensive exploration across diverse application domains remains necessary. Although we examined RQNN’s learning behavior through line rotation and ModelNet10 experiments, validation across various datasets and more complex object structures is essential. These investigations will help identify specific domains where RQNN’s unique characteristics, particularly the differences in generalization capabilities from existing models and its broader feature exploration in early learning stages, can be effectively leveraged for enhanced performance. The authors also feel that the work presented in this paper is probably only scratching the surface of what is possible by focusing on the inherent properties of hypercomplex algebras.
Acknowledgments.
The authors would like to thank the anonymous reviewers for valuable comments. This paper is part of the master’s thesis that Shogo Yamauchi, one of the authors, completed in 2024 at the Graduate School of Artificial Intelligence and Science, Rikkyo University.